OpenStudy (chosenmatt):

What is the value of the expression (–2)5? A. 32 B. 10 C. –10 D. –32 HELP ME

OpenStudy (anonymous):

C

OpenStudy (igreen):

\(-2 \times -2 \times -2 \times -2 \times -2 =~?\) @keynote Don't give answers.

sammixboo (sammixboo):

Is it (-2)5 or \((-2)^5\)?

sammixboo (sammixboo):

Yes, remember not to give out answers. Try to guide the asker to the answer :D

sammixboo (sammixboo):

@chosenmatt

OpenStudy (chosenmatt):

32?

OpenStudy (igreen):

Oh yeah.. ^^^^^^^^^^^^^^^^^

OpenStudy (igreen):

No, first. Is it \((-2)^5\) or is it \(-2(5)\)?

OpenStudy (chosenmatt):

(-2)5

OpenStudy (igreen):

Okay, are you sure the 5 isn't small and at the top?

OpenStudy (igreen):

If it isn't, then you just multiply: \(-2 \times 5\)

sammixboo (sammixboo):

Ah, okay, (-2)5 is the same as -2 x 5 When multiplying a negative with a positive, the answer is always negative.

OpenStudy (chosenmatt):

5 is at the top like 2 and an high 5 lol

OpenStudy (igreen):

Remember: \(Negative \times Positive = Negative\) \(Positive \times Negative = Negative\) \(Positive \times Positive = Positive\) \(Negative \times Negative = Negative\)

sammixboo (sammixboo):

So -2 and 5 is the same size?? o_o

OpenStudy (igreen):

Right, so it's: \(-2 \times -2 \times -2 \times -2 \times -2\)

OpenStudy (chosenmatt):

no 5 is on top of 2 lol

OpenStudy (igreen):

You mean -2, right?

OpenStudy (igreen):

Okay, if 5 is on the top of -2, like this: \((-2)^5\) Then you multiply -2 by itself, 5 times, giving you: \(-2 \times -2 \times -2 \times -2 \times -2\)

OpenStudy (chosenmatt):

yeah lol \[-2 ^{5}\]

OpenStudy (chosenmatt):

parentheses is around -2 tho

OpenStudy (chosenmatt):

@sleepyjess

OpenStudy (igreen):

Yeah, just multiply: \(-2 \times -2 \times -2 \times -2 \times -2\)

OpenStudy (chosenmatt):

s0 32?

OpenStudy (igreen):

No..check again.

OpenStudy (chosenmatt):

-32?

OpenStudy (igreen):

Yep! \(-2 \times -2 \times -2 \times -2 \times -2\) \(4 \times -2 \times -2 \times -2\) \(-8 \times -2 \times -2\) \(16 \times -2\) \(-32~\Large\color{lime} \checkmark\)